# Dominik Bojko Krzysztof Grining Marek Klonowski

### v6

@@ Appendix A Technical Lemmas and Proofs Related to Differential Privacy of Morris Counter
For the sake of completeness, we present here proofs of all technical lemmas that are not directly connected to Theorem 1. Some of computations are supported by Wolfram Mathematica ver.11.3 ([29]). Whenever we obtain a result in this manner, we indicate it by $\stackrel{{\scriptstyle W}}{{=}}$ sign. Usually results are precise, however in some cases, final forms are attained numerically.
@@
Next two lemmas will be useful in a proof of Lemma 6.
Lemma 4
@@
and $a(2cx)^{2y}-~a(cx)^{y}\leqslant~\dfrac{y}{4c^{2}x^{2}}~a(2cx)^{2(y-1)}$, which imply the thesis of this Lemma. $\qed$
Lemma 5
@@
For any possible $x\geqslant 8$ ($k\geqslant 3$), $\sum_{i=0}^{1}2^{-i(2i-1)}r_{2i+1}1.1^{1-i}L_{1}(x;u_{t}(i))>0.0015$. We already know that $W_{1}(i)$ are positive for $i>1$, so $p_{2^{k+1}+1,k+6}-p_{2^{k}+1,k+5}>0$ for all $k\geqslant 3$$\qed$
We may use Theorem 4 once again to see that ${p_{2^{6}+1,10}}/{p_{2^{6}+1,11}}=129.454\ldots>2^{7}$ and ${p_{2^{7}+1,11}}/{p_{2^{7}+1,12}}=125.065\ldots<2^{7}$. Together with Lemma 6 we may easily attain Claim 1 and we instantly see that this Claim cannot be extended continuously for $k<7$.
Lemma 7

### v7

@@ Appendix A Technical Lemmas and Proofs Related to Differential Privacy of Morris Counter
For the sake of completeness, we present here proofs of all technical lemmas that are not directly connected to Theorem 1. Some of computations are supported by Wolfram Mathematica ver.11.3 (Wolfram Research, Inc. (2025)). Whenever we obtain a result in this manner, we indicate it by $\stackrel{{\scriptstyle W}}{{=}}$ sign. Usually results are precise, however in some cases, final forms are attained numerically.
@@
Next two lemmas will be useful in a proof of Lemma A.21.
Lemma A.17.
@@
and $a(2cx)^{2y}-~a(cx)^{y}\leqslant~\dfrac{y}{4c^{2}x^{2}}~a(2cx)^{2(y-1)}$, which imply the thesis of this Lemma.
Lemma A.19.
@@
For any possible $x\geqslant 8$ ($k\geqslant 3$), $\sum_{i=0}^{1}2^{-i(2i-1)}r_{2i+1}1.1^{1-i}L_{1}(x;u_{t}(i))>0.0015$. We already know that $W_{1}(i)$ are positive for $i>1$, so $p_{2^{k+1}+1,k+6}-p_{2^{k}+1,k+5}>0$ for all $k\geqslant 3$.
We may use Theorem 4 once again to see that
$\frac{p_{2^{6}+1,10}}{p_{2^{6}+1,11}}=129.454\ldots>2^{7}\mbox{and},$
$\frac{p_{2^{7}+1,11}}{p_{2^{7}+1,12}}=125.065\ldots<2^{7}.$
Together with Lemma A.21 we may easily attain Claim 1 and we instantly see that this Claim cannot be extended continuously for $k<7$.
Lemma A.23.